Can someone help me with this question, I do not know how to do the long-distance in excel.
If you could show me the Excel data as well that would be great.
retention 10.18 The following are the average distances of the planets in the solar system from the 25 Distance (millions of miles) 47.163 67.235 92.960 141.61 313.00 483.60 886.70 1783.0 2794.0 3666.0 Planet No. Planet Mercury Venus Earth Mars Asteroids Jupiter Saturn Uranus Neptune Pluto 4 10 (a) How does the distance of a planet from the sun increase with the planet number? Find a transformation of the distance that gives a linear relationship with respect to the planet number. (b) Fit a least squares straight line after linearizing the relationship. (c) It is speculated that there is a planet beyond Pluto, called Planet X. Predict its distance from the sun.
Answer #1

Hello,
Here the planet number is the depend on the distance.
we fit the linear regression model.

Y=a+bx+e

Where y- dependent variable

X-independent variable

A-intercept

b – Regression coefficient

E-random error

Input data in theExcel,
follow the path
data>datat analysis >regression
the excel output of the above regression is

Regression Statistics

Multiple R

0.887914

R Square

0.788392

Adjusted R Square

0.761941

Standard Error

630.3226

Observations

10

Coefficients

Standard Error

t Stat

P-value

Lower 95%

Upper 95%

Lower 95.0%

Upper 95.0%

Intercept

-1056.24

430.5923

-2.45299

0.039751

-2049.19

-63.2909

-2049.19

-63.2909

planet no

378.8664

69.39623

5.459467

0.000602

218.8384

538.8944

218.8384

538.8944

Here planet number is significant and r-square is 0.78.

PREDICTION OF y,

The estimated regression equation as,
y=-1056.24+378.86X   ……………estimated regression equation
if the x=x then regression equation as

y=-1056.24+378.86X   , after substituting the value of the x we can get the distance.
thanks